# Large Language Models Are State-of-the-Art Evaluator for Grammatical Error Correction

### v1

@@ Abstract
Large Language Models (LLMs) have been reported to outperform existing automatic evaluation metrics in some tasks, such as text summarization and machine translation. However, there has been a lack of research on LLMs as evaluators in grammatical error correction (GEC). In this study, we investigate the performance of LLMs in GEC evaluation by employing prompts designed to incorporate various evaluation criteria inspired by previous research. Our extensive experimental results demonstrate that GPT-4 achieved Kendall’s rank correlation of 0.662 with human judgments, surpassing all existing methods. Furthermore, in recent GEC evaluations, we have underscored the significance of the LLMs scale and particularly emphasized the importance of fluency among evaluation criteria.
@@ 1 Introduction
In recent years, several studies have been conducted on the use of LLMs as an evaluator. In text summarization, dialogue generation, and machine translation, GPT-4 has demonstrated superior performance compared to existing automatic evaluation metrics Liu et al. (2023b); Kocmi and Federmann (2023). While there is currently no specific research on GEC evaluation, considering GPT-4’s ability to explain grammatical errors with 90% accuracy in human evaluations Song et al. (2023), it holds potential for evaluating corrections.
@@
Therefore, we aim to explore the extent to which LLMs operate as evaluation models in GEC. Specifically, we conduct GEC evaluations using LLMs with prompts at different evaluation granularities to investigate how evaluation capabilities change with the presence of evaluation criteria and the scale of LLMs, as shown in Figure 1. Kobayashi et al. (2024)’s work on the evaluation of metrics (i.e., meta-evaluation) has revealed that conventional metrics lack the resolution to capture performance differences in high-performing GEC systems. Given this current state, to facilitate proper GEC evaluation moving forward, we investigate the potential of LLMs by comparing them with conventional metrics through meta-evaluation.
Our contributions are summarized as follows. (1) We conducted the first investigation into the performance of LLMs as evaluators in GEC for the first time, and the results showed that GPT-4 achieved state-of-the-art performance, indicating the usefulness of considering evaluation criteria in prompts (especially fluency). (2) It was suggested that as LLM scales decrease, the correlation with human judgments decreases, and the ability to capture fluency in corrected sentences diminishes. Smaller LLMs tend to avoid extreme scores, while larger LLMs tend to assign higher scores.
@@ 2.1 Considered metrics
For EBMs, we employ four GEC metrics. $M^{2}$ Dahlmeier and Ng (2012) dynamically extracts edits to maximize overlap with gold annotations from the corrected text and calculates the F-score. ERRANT Bryant et al. (2017) is similar to $M^{2}$, but distinct in utilizing a linguistically enhanced alignment algorithm for edit extraction. GoToScorer Gotou et al. (2020) calculates an F-score taking into account the difficulty of corrections. PT-$M^{2}$ Gong et al. (2022) combines $M^{2}$ with BERTScore Zhang et al. (2019), enabling the measurement of semantic similarity in addition to simply comparing edits.
For SBMs, we utilize four GEC metrics. GLEU Napoles et al. (2015) rewards $n$-grams in the corrected sentence that match the reference but are not in the source, while penalizing $n$-grams in the source that do not match the reference. We use GLEU without tuning Napoles et al. (2016). Scribendi Score Islam and Magnani (2021) evaluates based on GPT-2 perplexity, token sort ratio, and Levenshtein distance ratio. SOME Yoshimura et al. (2020) fine-tunes BERT Devlin et al. (2019) using human evaluation scores based on three criteria: grammaticality, fluency, and semantic preservation. IMPARA Maeda et al. (2022) utilizes a quality estimation model and a similarity model based on BERT to consider the impact of edits.
@@
2.2 Meta-evaluation methods
We conduct system-level and sentence-level meta-evaluations using SEEDA dataset Kobayashi et al. (2024). SEEDA consists of human ratings at two different granularities: edit-based and sentence-based, for 12 modern GEC systems and human-authored sentences. The dataset comprises two components: SEEDA-E based on edit-based evaluation and SEEDA-S based on sentence-based evaluation. We conduct two variations of meta-evaluation: “Base”, which uses the 12 systems excluding outliers, and “+ Fluent corr.”, which adds two fluent corrected sentences [^1] In GEC, there are two types of edits: minimal edits, which make the minimum necessary corrections, and fluency edits, which aim to make the sentence more fluent. additionally.
System-level meta-evaluation:
In the system-level meta-evaluation, we utilize the system scores derived from human rankings using TrueSkill Sakaguchi et al. (2014). For metrics like SOME, where system-level scores cannot be directly calculated, we use the average of sentence-level scores as a substitute. Additionally, for LLMs, we employ system scores derived from LLMs rankings (Appendix B) similar to human rankings. To measure the correlation between human judgments and metric scores, we use Pearson correlation ($r$) and Spearman rank correlation ($\rho$). To ensure proper correlation calculation, we use the set of sentences that humans evaluated to compute the metric scores.
Sentence-level meta-evaluation:
In the sentence-level meta-evaluation, we use pairwise judgments from SEEDA. To investigate the proximity between human judgments and metric scores, we employ Accuracy (Acc) and Kendall’s rank correlation ($\tau$). Kendall ($\tau$) is valuable for assessing performance in common use cases where corrections are compared to each other.
@@ 2.2 Meta-evaluation methods
Figure 2: Window analysis was performed by selecting any consecutive four systems from the human rankings of the 12 systems (“Base”). For instance, x=5 involves calculating the Pearson correlation ($r$) using the systems ranked from 5th to 2nd in the human rankings. In contrast to conventional GEC metrics, which exhibit unstable correlations, GPT-4 demonstrates relatively stable correlations.
@@ 3.2 Sentence-level analysis
In the sentence-level meta-evaluation, we observed differences in correlations between metrics that were not apparent in the system-level meta-evaluation. In particular, while GPT-4-E and GPT-4-S showed similar correlations in system-level meta-evaluation, it was revealed that there was a significant difference between them. Additionally, considering fluent corrected sentences (“+ Fluent corr.”) led to a slight decrease in overall correlation, but GPT-4 still maintained a considerably high correlation compared to traditional metrics. This suggests that GPT-4 exhibits strong correlations with human judgments and that examining sentence-level correlations is beneficial for comparing high-performance metrics.
@@ 3.3 Further analysis
To increase the difficulty of the meta-evaluation task, we computed correlations using a set of systems with similar performance. According to the window analysis in Figure 3, GPT-4 maintains relatively high and stable correlations, making them suitable for evaluating modern neural systems in recent years. In SEEDA-E, the notably high correlations of GPT-4-S + Fluency across almost all data points emphasize the importance of fluency. In SEEDA-S, while overall correlations are high, the significant decrease in correlation at x=10, suggests the presence of GEC systems that are challenging to evaluate for the metrics. On the other hand, conventional metrics frequently exhibit either no correlation or negative correlation, indicating their low robustness in GEC evaluation.
@@ 5 Conclusion
In this work, we investigated the capability of LLMs as evaluators in GEC, and GPT-4 demonstrated significantly higher correlations compared to traditional metrics. Future work should delve into the impact of few-shot learning and optimize prompt engineering for enhanced evaluation performance. Furthermore, we plan to explore the possibility of document-level evaluation, considering the expansion of the GPT’s context window, which is not currently focused on by existing metrics.
@@ Appendix C Tendency of LLM scoring
The distribution of scores assigned by LLMs to corrected sentences is shown in Figure 4. As the LLM scale increases, there is a tendency to assign higher scores (4 or 5 points). Based on our meta-evaluation results, which suggest that higher LLM scales are associated with higher correlations with human judgments, smaller LLMs may underestimate corrections judged to be good by humans. Llama 2-S tends to avoid extreme scores such as 1 or 5 points and shows a high degree of score overlap, making it difficult to compare more detailed corrected sentences.

### v2

@@ Abstract
Large Language Models (LLMs) have been reported to outperform existing automatic evaluation metrics in some tasks, such as text summarization and machine translation. However, there has been a lack of research on LLMs as evaluators in grammatical error correction (GEC). In this study, we investigate the performance of LLMs in English GEC evaluation by employing prompts designed to incorporate various evaluation criteria inspired by previous research. Our extensive experimental results demonstrate that GPT-4 achieved Kendall’s rank correlation of 0.662 with human evaluations, surpassing all existing methods. Furthermore, in recent GEC evaluations, we have underscored the significance of the LLMs scale and particularly emphasized the importance of fluency among evaluation criteria.
@@ 1 Introduction
In recent years, several studies have been conducted on the use of LLMs as an evaluator. In text summarization, dialogue generation, and machine translation, GPT-4 has demonstrated superior performance compared to existing automatic evaluation metrics Liu et al. (2023b); Kocmi and Federmann (2023). While there is very little research on GEC evaluation, considering GPT-4’s ability to explain grammatical errors with 90% accuracy in human evaluations Song et al. (2023), it holds potential for evaluating corrections. Sottana et al. (2023) conducted meta-evaluation using a limited number of systems, but there has been no comprehensive analysis using dozens of systems like traditional approaches such as Grundkiewicz et al. (2015) and Kobayashi et al. (2024).
@@
Therefore, we aim to explore the extent to which LLMs operate as evaluation models in English GEC. Specifically, we conduct GEC evaluations using LLMs with prompts at different evaluation granularities to investigate how evaluation capabilities change with the presence of evaluation criteria and the scale of LLMs, as shown in Figure 1. Kobayashi et al. (2024)’s work on the evaluation of metrics (i.e., meta-evaluation) has revealed that conventional metrics lack the resolution to capture performance differences in high-performing GEC systems. Given this current state, to facilitate proper GEC evaluation moving forward, we investigate the potential of LLMs by comparing them with conventional metrics through meta-evaluation.
Our contributions are summarized as follows. (1) We conducted a comprehensive investigation into the performance of LLMs as evaluators in GEC, and the results showed that GPT-4 achieved state-of-the-art performance, indicating the usefulness of considering evaluation criteria in prompts (especially fluency). (2) It was suggested that as LLM scales decrease, the correlation with human evaluations decreases, and the ability to capture fluency in corrected sentences diminishes. Smaller LLMs tend to avoid extreme scores, while larger LLMs tend to assign higher scores.
@@ 2.1 Considered metrics
For EBMs, we employ four metrics.
  • $M^{2}$ Dahlmeier and Ng (2012) dynamically extracts edits using Levenshtein algorithm to maximize overlap with gold annotations from the hypothesis sentences and calculates the F-score.
  • ERRANT Bryant et al. (2017) is similar to $M^{2}$, but it differs in that it uses a linguistically extended Damerau-Levenshtein algorithm for edit extraction to enhance the alignment of tokens with similar linguistic properties.
  • GoToScorer Gotou et al. (2020) calculates an F-score taking into account the difficulty of corrections. The difficulty is defined based on the number of systems that could correctly correct errors per total number of systems.
  • PT-$M^{2}$ Gong et al. (2022) combines $M^{2}$ with BERTScore Zhang et al. (2019), enabling the measurement of semantic similarity in addition to simply comparing edits.
For SBMs, we utilize four metrics.
  • GLEU Napoles et al. (2015) rewards $n$-grams in the hypothesis sentence that match the reference but are not in the source sentence while penalizing $n$-grams in the source that do not match the reference. We use GLEU without tuning Napoles et al. (2016).
  • Scribendi Score Islam and Magnani (2021) evaluates based on GPT-2 perplexity, token sort ratio, and Levenshtein distance ratio.
  • SOME Yoshimura et al. (2020) fine-tunes BERT Devlin et al. (2019) using human evaluation scores based on three criteria: grammaticality, fluency, and meaning preservation.
  • IMPARA Maeda et al. (2022) utilizes a quality estimation model and a similarity model based on BERT to consider the impact of edits.
@@
2.2 Meta-evaluation methods
We conduct system-level and sentence-level meta-evaluations using SEEDA dataset Kobayashi et al. (2024). SEEDA consists of human evaluations at two different granularities: edit-based and sentence-based, for 12 outputs from neural-based GEC systems and 3 human-authored sentences. The dataset comprises two components: SEEDA-E based on edit-based evaluation and SEEDA-S based on sentence-based evaluation. In SEEDA, for correction pairs (A, B) sampled from these corrected sentence collections, three annotators provide 5-point scores for each granularity, resulting in 5347 pairwise judgments (A$>$B, A$=$B, A$<$B). Subsequently, human rankings (from 1st to 15th place) of systems are obtained from pairwise judgments using rating algorithms such as Trueskill Sakaguchi et al. (2014) and Expected Wins Bojar et al. (2013). We conduct two variations of meta-evaluation: “Base”, which uses the 12 systems excluding outliers, and “+ Fluent corr.”, which adds two fluent corrected sentences [^1] In GEC, there are two types of edits: minimal edits, which make the minimum necessary corrections, and fluency edits, which aim to make the sentence more fluent. additionally.
System-level meta-evaluation:
In the system-level meta-evaluation, we utilize the system scores derived from human rankings of systems using TrueSkill Sakaguchi et al. (2014). For metrics like SOME, where system-level scores cannot be directly calculated, we use the average of sentence-level scores as a substitute. Additionally, for LLMs, we employ system scores derived from LLMs rankings (Appendix B) similar to human rankings. To measure the correlation between human evaluations and metric scores, we use Pearson correlation ($r$) and Spearman rank correlation ($\rho$). To ensure proper correlation calculation, we use the set of sentences that humans evaluated to compute the metric scores.
Sentence-level meta-evaluation:
In the sentence-level meta-evaluation, we use pairwise judgments from SEEDA. To investigate the proximity between human evaluations and metric scores, we employ Accuracy (Acc) and Kendall’s rank correlation ($\tau$). Kendall ($\tau$) is valuable for assessing performance in common use cases where corrections are compared to each other.
Figure 2: Window analysis was performed by selecting any consecutive four systems from the human rankings of the 12 systems (“Base”). For instance, x=4 involves calculating the Pearson correlation ($r$) using the systems ranked from 1st to 4th in the human rankings. In contrast to conventional GEC metrics, which exhibit unstable correlations, GPT-4 demonstrates relatively stable correlations.
@@ 3.2 Sentence-level analysis
In the sentence-level meta-evaluation, we observed differences in correlations between metrics that were not apparent in the system-level meta-evaluation. In particular, while GPT-4-E and GPT-4-S showed similar correlations in system-level meta-evaluation, it was revealed that there was a notable difference between them. Additionally, considering fluent corrected sentences (“+ Fluent corr.”) led to a slight decrease in overall correlation, but GPT-4 still maintained a considerably high correlation compared to traditional metrics. This suggests that GPT-4 exhibits strong correlations with human evaluations and that examining sentence-level correlations is beneficial for comparing high-performance metrics.
@@ 3.3 Further analysis
To increase the difficulty of the meta-evaluation task, we computed correlations using a set of systems with similar performance. Specifically, we conduct system-level meta-evaluation using only subsets of consecutive four systems in the human rankings of systems, and show the transitions of correlation at positions from 4th to 12th as window analysis in Figure 2333For simplicity, we exclude the results of Llama 2 and GPT-3.5, which showed low performance.. For example, the point at x=4 represents the Pearson correlation value calculated using only the outputs of the four systems ranked from 1st to 4th.
According to the window analysis, GPT-4 maintains relatively high and stable correlations, making them suitable for evaluating modern neural systems in recent years. In SEEDA-E, the notably high correlations of GPT-4-S + Fluency across almost all data points emphasize the importance of fluency. In SEEDA-S, while overall correlations are high, the significant decrease in correlation at x=10, suggests the presence of GEC systems that are challenging to evaluate for the metrics. On the other hand, conventional metrics frequently exhibit either no correlation or negative correlation, indicating their low robustness in GEC evaluation.
@@ 5 Conclusion
In this work, we investigated the capability of LLMs as evaluators in English GEC, and GPT-4 demonstrated significantly higher correlations compared to traditional metrics. Future work should delve into the impact of few-shot learning and optimize prompt engineering for enhanced evaluation performance. Furthermore, we plan to explore the possibility of document-level evaluation, considering the expansion of the GPT’s context window, which is not currently focused on by existing metrics.
@@ 6 Limitations
Acknowledgments
This work is partly supported by JST, PRESTO Grant Number JPMJPR2366, Japan.
@@ Appendix C Tendency of LLM scoring
The distribution of scores assigned by LLMs to corrected sentences is shown in Figure 4. As the LLM scale increases, there is a tendency to assign higher scores (4 or 5 points). Based on our meta-evaluation results, which suggest that higher LLM scales are associated with higher correlations with human evaluations, smaller LLMs may underestimate corrections judged to be good by humans. Llama 2-S tends to avoid extreme scores such as 1 or 5 points and shows a high degree of score overlap, making it difficult to compare more detailed corrected sentences.